Switch Problem -15 12R Ele opened at a certain time. much does it take tie ouvrent Mico of inital value
Answer #1

During discharging the current decays as i = I*e^-(t/RC)

I = iniital current

C = capacitance = 3*10^-15 F

i = I/100


I/100 = I*e^-(t/(12*3*10^-15))


1/100 = e^-(t/(12*3*10^-15))

t = 1.66*10^-13 s   <<-------answer